OpenStudy (migitmack):

What are the solutions of x2 + 8x − 20 = 0? Please show your work

OpenStudy (anonymous):

x^2+8x-20=0 +20 +20 x^2+8x=20 x^2=1 soo... 1x+8x=9x 9x=20 now u divide 20/9

OpenStudy (campbell_st):

factorise the quadratic find the factors of -20 that add to 8 10 and -2 (x +10)(x - 2) = 0 to solve you need to solve the equations x + 10 = 0 x-2 = 0 so the solutions to the equation are x = -10 or 2